OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -2 f(1) = 8 f(n) = -4 f(n -1) - 3 f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? (4 points) -2, 8, -26, -80, -2, 8, -26, 80, -2, 12, -44, 180, -2, -12, -44, -180,

OpenStudy (anonymous):

\[f(2)=-4\times f(1)-3\times f(0)\\ f(2)=-4\times 8-3\times (-2)\]

OpenStudy (anonymous):

What am i suppose to do next ?